OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ x ^{2}-3x-18 }{ x ^{2}-36 }\]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

x-6/x-3

OpenStudy (anonymous):

x+6/x-6

OpenStudy (anonymous):

x-3/x+3

OpenStudy (anonymous):

x+3/x+6

OpenStudy (skullpatrol):

Do you recognise a difference of two squares in the denominator?

OpenStudy (anonymous):

36 divided be 18 is 2

OpenStudy (anonymous):

by

OpenStudy (skullpatrol):

$$\huge x^2-36=x^2-6^2$$

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Recall $$\huge a^2-b^2=?$$

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

$$\huge a^2-b^2=(a+b)(a-b)$$

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

ok so??

OpenStudy (akashdeepdeb):

skullpatrol will help you.

OpenStudy (skullpatrol):

Please apply this formula to the denominator of your question.

OpenStudy (skullpatrol):

I can help you learn, but I can't learn for you :-)

OpenStudy (anonymous):

k but it makes no sense

OpenStudy (skullpatrol):

What makes no sense?

OpenStudy (skullpatrol):

Apply this formula: $$a^2-b^2=(a+b)(a-b)$$ to $$x^2-6^2$$

OpenStudy (skullpatrol):

Please try to learn...

OpenStudy (skullpatrol):

In 2 minutes I'm gone.

OpenStudy (skullpatrol):